# DVMNet: Computing Relative Pose for Unseen Objects Beyond Hypotheses

### v1

@@ Abstract
Determining the relative pose of an object between two images is pivotal to the success of generalizable object pose estimation. Existing approaches typically approximate the continuous pose representation with a large number of discrete pose hypotheses, which incurs a computationally expensive process of scoring each hypothesis at test time. By contrast, we present a Deep Voxel Matching Network (DVMNet) that eliminates the need for pose hypotheses and computes the relative object pose in a single pass. To this end, we map the two input RGB images, reference and query, to their respective voxelized 3D representations. We then pass the resulting voxels through a pose estimation module, where the voxels are aligned and the pose is computed in an end-to-end fashion by solving a least-squares problem. To enhance robustness, we introduce a weighted closest voxel algorithm capable of mitigating the impact of noisy voxels. We conduct extensive experiments on the CO3D, LINEMOD, and Objaverse datasets, demonstrating that our method delivers more accurate relative pose estimates for novel objects at a lower computational cost compared to state-of-the-art methods. Our code is released at: https://github.com/sailor-z/DVMNet/.
@@ 1 Introduction
Object pose estimation plays a crucial role in 3D computer vision and robotics tasks [2, 11, 23, 33], aiming to estimate the 3D rotation and 3D translation of an object depicted in an RGB image. The vast majority of existing methods work under the assumption that the training and testing data include the same object instances, thereby limiting their applicability to scenarios that involve previously unseen objects. Recently, generalizable object pose estimation [21, 29, 32] has received growing attention, showcasing the potential to generalize to unseen objects from new categories without retraining the network. In pursuit of this generalization capability, existing methods leverage densely sampled images depicting unseen objects in diverse poses, serving as references. Object pose estimation is then carried out through template matching [21, 29] or by establishing 2D-3D correspondences [32, 14]. Unfortunately, the effectiveness of these methods strongly depends on the references densely covering the viewpoints of unseen objects, making them inapplicable to practical scenarios where only sparse reference views are available.
@@
Figure 1: Advantages of our DVMNet compared to hypothesis-based methods. Hypothesis-based techniques approximate the relative object pose by scoring numerous pose hypotheses, leading to a high computational cost. By contrast, our DVMNet computes the pose in a hypothesis-free fashion by robustly matching voxelized 3D representations of the reference and query images via a Weighted Closest Voxel (WCV) algorithm. Our method strikes a favorable balance between computational cost and accuracy in relative object pose estimation, as measured by multiply-accumulate operations (MACs) and angular error.
@@
In this context, a few methods [47, 18, 52] highlight the importance of relative object pose estimation. Unlike previous approaches in generalizable object pose estimation, these methods focus on estimating the relative pose of an unseen object between two images, i.e., a single reference image of the object and the query image. In this paper, we also work in this setting, motivated by the practical ease of obtaining a single reference image for a new object. One plausible solution is to compute the relative pose based on 2D-2D correspondences [13]. However, the single-reference scenario tends to yield a significant viewpoint gap between the reference and the query. Existing studies [18, 52] have shown that image-matching techniques [26, 31] are sensitive to such pose differences. To handle this issue, the prior methods [47, 18, 52] follow an alternative strategy of scoring multiple pose hypotheses for the input reference-query pair, and predicting the pose as the hypothesis with the highest score. However, this strategy comes with the drawback of requiring numerous pose hypotheses to achieve reasonable accuracy, e.g., 500,000 in [18], which thus induces a computational burden. Moreover, we empirically found that these approaches occasionally produce unnaturally large errors. One plausible explanation is their failure to model the continuous nature of the object pose space, as they primarily concentrate on learning to score discrete hypotheses.
@@
To overcome these drawbacks, we present a novel Deep Voxel Matching Network (DVMNet) that computes the relative pose of unseen objects in a hypothesis-free and end-to-end manner. Our approach starts by voxelizing the query and reference images in a dedicated autoencoder. The encoder network lifts 2D image features to 3D voxels, leveraging cross-view 3D information. The decoder network reconstructs a masked object image from the voxels, encouraging the learned voxels to account for the object. Subsequently, we align the query and reference voxels based on a score matrix that measures the voxel similarities. To handle unreliable voxels due to background, varying illumination, and other potential nuisances, we present a Weighted Closest Voxel (WCV) algorithm to facilitate robust pose estimation. In this algorithm, each voxel-voxel correspondence is assigned a confidence score computed by utilizing both the 3D voxel objectness map and the 2D object mask learned by the autoencoder. The relative object pose is then computed by solving a weighted least-squares problem. Such an end-to-end learning mechanism eliminates the necessity for voxel-wise annotations and allows the network to directly learn pose-aware features from RGB images. As illustrated in Fig. 1, our DVMNet requires significantly fewer multiply-accumulate operations (MACs) while achieving smaller angular errors than its hypothesis-based competitors.
@@
We perform comprehensive experiments on the CO3D [25], LINEMOD [15], and Objaverse [9] datasets. Our method yields more accurate and robust relative pose estimates for previously unseen objects than both image-matching and hypothesis-based competitors. We also conduct ablation studies where the results demonstrate the effectiveness of the key components in our framework. In short, our contributions are twofolds: First, we tackle the problem of relative pose estimation for unseen objects in a hypothesis-free manner by introducing a deep voxel matching network. Second, we present a weighted closest voxel algorithm that robustly and end-to-end computes the object pose from voxel-voxel correspondences.
@@ 2 Related Work
Instance-Level Object Pose Estimation. The majority of previous deep learning approaches to object pose estimation [43, 24, 35, 30, 34] tackle the problem at an instance level, holding an assumption that the training and testing data depict the same object instances. Since the appearance of an object instance in different poses typically exhibits limited variations, these methods provide highly accurate object pose estimates. Nevertheless, they struggle to generalize to previously unseen objects during testing without retraining the network, as has been observed in the literature [21, 29, 32]. This limitation constrains their applicability in real-world scenarios that often involve diverse object instances. This has been remedied to a degree by category-level object pose estimation methods [36, 7, 19]. In this scenario, the testing images comprise new object instances from specific categories already included in the training data. Although these methods have achieved promising generalization ability within the predefined object categories, they become ineffective when facing objects from entirely new categories.
@@
Generalizable Object Pose Estimation. To tackle the scenario of unseen objects from new categories, there has been growing interest in generalizable object pose estimation. When a textured 3D mesh is available for an unseen object, some approaches [42, 51, 29] suggest generating synthetic images as references by rendering the 3D mesh from various viewpoints. Given a query image that depicts this object, a template matching paradigm is utilized to identify the most similar reference and approximate the object pose in the query as that of the selected reference. Some methods bypass the need for 3D meshes by assuming the availability of multiple real reference images. Object pose estimation is then carried out by employing either a template matching strategy [21] or a 3D object reconstruction technique [32, 14]. Nevertheless, all of these methods rely on having access to dense-view reference images, which limits their applicability in scenarios where only sparse views of references are available.
@@
Relative Object Pose Estimation. Recently, several studies [47, 18, 52, 37] have highlighted the importance of relative object pose estimation. These methods stand out in generalizable object pose estimation due to their key advantage of requiring only a single reference image. The objective of these methods is to estimate the relative object pose between the input query image and the reference. Since the single-reference assumption tends to result in a large object pose difference between the query and the reference, unseen object pose estimation becomes more challenging. Intuitively, one could establish pixel-pixel correspondences between the two images and compute the relative object pose based on multi-view geometry [13]. However, as reported in the literature [18, 52] and also in our experiments, even the state-of-the-art image-matching techniques [26, 31, 12] cannot deliver accurate pose estimates when confronted with large object pose differences. To address this, existing methods [47, 18, 52] suggest approximating the relative object pose via a discrete set of pose hypotheses, and learning to maximize scores of the positive hypotheses. Since object pose lies in a continuous space [53], accurately approximating the pose necessitates a vast number of pose hypotheses, which makes the hypothesis-based approaches computationally expensive. Moreover, scoring discrete samples lacks an understanding of the continuous pose distribution, leading to failure cases with unnaturally high pose estimation errors. By contrast, we present a hypothesis-free technique that is capable of computing the relative object pose in a single pass via deep voxel matching.
@@ 3 Method
Figure 2: Network architecture of our autoencoder. The encoder takes two RGB images, query and reference, as input and lifts their 2D feature embeddings to 3D voxels by leveraging cross-view 3D information. $\mathbf{O}_{q}$ and $\mathbf{O}_{r}$ represent the learned 3D objectness maps account for robust object pose estimation. The decoder then reconstructs the masked object images from the voxels, allowing the voxels to encode the object patterns.
@@ 3.1 Problem Formulation
We tackle the problem of estimating the relative pose $\Delta{\mathbf{P}}$ of a previously-unseen object between a query image $\mathbf{I}_{q}$ and a reference image $\mathbf{I}_{r}$. In this scenario, the objects present in the testing set $\Omega_{\text{test}}$ differ from those in the training set $\Omega_{\text{train}}$, and the goal is to handle the unseen objects without retraining the network. This requires addressing the challenges of generalization to unseen objects in $\Omega_{\text{test}}$ and robustness to the large object pose difference between $\mathbf{I}_{q}$ and $\mathbf{I}_{r}$.
Since the relative object translation estimation can be facilitated by using zero-shot object detectors [50, 39, 17], our focus is on predicting the relative object rotation, i.e., $\Delta{\mathbf{R}}\in{SO(3)}$, which is more challenging. This choice aligns with existing literature [42, 44, 51, 52]. Previous hypothesis-based approaches [47, 18, 52] approximate $\Delta{\mathbf{R}}$ by sampling discrete pose hypotheses and maximizing the scores of positive samples. The problem is formulated as
@@
where $\mathcal{R}$ denotes the set of discrete pose hypotheses. Achieving a decent approximation accuracy requires a large number of hypotheses, e.g., 500,000 in [18]. By contrast, we present a hypothesis-free technique that computes $\Delta{\mathbf{R}}$ in a single pass as $\Delta{\hat{\mathbf{R}}}=g(\mathbf{I}_{q},\mathbf{I}_{r})$.
3.2 Motivation

### v2

# DVMNet++: Rethinking Relative Pose Estimation for Unseen Objects

@@ Abstract
Determining the relative pose of a previously unseen object between two images is pivotal to the success of generalizable object pose estimation. Existing approaches typically predict 3D translation utilizing the ground-truth object bounding box and approximate 3D rotation with a large number of discrete hypotheses. This strategy makes unrealistic assumptions about the availability of ground truth and incurs a computationally expensive process of scoring each hypothesis at test time. By contrast, we rethink the problem of relative pose estimation for unseen objects by presenting a Deep Voxel Matching Network (DVMNet++). Our method computes the relative object pose in a single pass, eliminating the need for ground-truth object bounding boxes and rotation hypotheses. We achieve open-set object detection by leveraging image feature embedding and natural language understanding as reference. The detection result is then employed to approximate the translation parameters and crop the object from the query image. For rotation estimation, we map the two RGB images, i.e., reference and cropped query, to their respective voxelized 3D representations. The resulting voxels are passed through a rotation estimation module, which aligns the voxels and computes the rotation in an end-to-end fashion by solving a least-squares problem. To enhance robustness, we introduce a weighted closest voxel algorithm capable of mitigating the impact of noisy voxels. We conduct extensive experiments on the CO3D, Objaverse, LINEMOD, and LINEMOD-O datasets, demonstrating that our approach delivers more accurate relative pose estimates for novel objects at a lower computational cost compared to state-of-the-art methods. Our code is released at:https://github.com/sailor-z/DVMNet/.
Index Terms:
Object pose estimation, unseen objects, two-view geometry, 3D computer vision.
@@ 1 Introduction
Object pose estimation plays a crucial role in 3D computer vision and robotics tasks [1, 2, 3, 4], aiming to produce 3D translation and 3D rotation of an object depicted in an RGB image. The vast majority of existing methods work under the assumption that the training and testing data include the same object instances, thereby limiting their applicability to scenarios that involve previously unseen objects. Recently, generalizable object pose estimation [5, 6, 7, 8] has received growing attention, showcasing the potential to generalize to unseen objects from new categories without retraining the network. In pursuit of this generalization capability, existing methods leverage densely sampled images depicting unseen objects in diverse poses, serving as references. Object pose estimation is then carried out through template matching [5, 6, 9, 8] or by establishing 2D-3D correspondences [7, 10, 11]. Unfortunately, the effectiveness of these methods strongly depends on the references densely covering the viewpoints of the unseen objects, making them inapplicable to practical scenarios where only sparse reference views are available.
@@
Figure 1: Advantages of our DVMNet++ compared to hypothesis-based methods. Hypothesis-based techniques approximate the relative object rotation by scoring numerous rotation hypotheses, leading to a high computational cost. By contrast, our DVMNet++ computes the rotation in a hypothesis-free fashion by robustly matching voxelized 3D representations of the reference and query images via a Weighted Closest Voxel algorithm. Our method strikes a favorable balance between computational cost and accuracy in relative object pose estimation, as measured by multiply-accumulate operations (MACs) and angular error.
@@
In this context, a few methods [12, 13, 14] highlight the importance of relative object pose estimation. Unlike previous approaches in generalizable object pose estimation, these methods focus on estimating the relative pose of an unseen object between two images, i.e., a query image and a single reference image of the object. In this paper, we also work in this setting, motivated by the practical ease of obtaining a single reference image for a new object. One plausible solution is to compute the relative pose based on 2D-2D correspondences [15]. However, the single-reference scenario tends to yield a significant viewpoint gap between the reference and the query. Existing studies [13, 14] have shown that image-matching techniques [16, 17] are sensitive to such pose differences. To handle this issue, the prior methods [12, 13, 14] follow an alternative strategy of scoring multiple rotation hypotheses for the input reference-query pair, and predicting the rotation as the hypothesis with the highest score. However, this alternative comes with the drawback of requiring numerous rotation hypotheses to achieve reasonable accuracy, e.g., 500,000 in [13], which thus induces a computational burden. Moreover, we empirically found that these approaches occasionally produce unnaturally large errors. One plausible explanation is their failure to model the continuous nature of the object rotation space, as they primarily concentrate on learning to score discrete hypotheses.
Additionally, it is worth noting that the aforementioned approaches assume that the ground-truth object bounding boxes are known, even at test time. Such ground-truth information facilitates the relative object pose estimation from two aspects: First, the object bounding box parameters are fed into a translation regression network [12, 13], which provides strong prior information about the object translation; second, the region containing the object is cropped from the query image, mitigating the impact of the background when estimating the object rotation. Unfortunately, since we focus on relative pose estimation for novel objects that are not included in the training set, detecting the previously unseen objects is non-trivial, particularly in cluttered scenes [18]. Due to the reliance on the ground-truth object bounding boxes, existing methods become inapplicable in scenarios where high-accuracy object bounding boxes are unavailable.
@@
To overcome these drawbacks, we present a new pipeline DVMNet++ that computes the relative pose of unseen objects efficiently without relying on ground-truth object bounding boxes. Our approach starts by detecting the object in the query image. Specifically, we draw inspiration from recent progress in open-vocabulary object detection [19, 20, 21, 22], which demonstrates promising detection accuracy for unseen objects. Since we have access to the reference image depicting the object, we describe the object via text prompts based on the reference image. The text prompts and the query image are taken as the input of an open-vocabulary object detection network [21] that produces object proposals. To identify the most reliable object bounding box from the generated proposals, we measure the similarity of the reference image and each proposal in high-dimensional feature space. The proposal closest to the reference in the feature space is selected as the detection result. We approximate the relative object translation and crop the object from the query using the identified bounding box. Subsequently, we achieve the hypothesis-free relative object rotation estimation by introducing a deep voxel matching network. We first voxelize the reference image and the cropped query image in a dedicated autoencoder. The encoder network lifts 2D image features to 3D voxels, leveraging cross-view 3D information. The decoder network reconstructs a masked object image from the voxels, encouraging the learned voxels to account for the object. We then align the query and reference voxels based on a score matrix that measures the voxel similarities. To handle unreliable voxels due to background, varying illumination, and other potential nuisances, we present a Weighted Closest Voxel (WCV) algorithm to facilitate robust rotation estimation. In this algorithm, each voxel-voxel correspondence is assigned a confidence score computed by utilizing both the 3D voxel objectness map and the 2D object mask learned by the autoencoder. The relative object rotation is computed by solving a weighted least-squares problem. Such an end-to-end learning mechanism eliminates the necessity for voxel-wise annotations and allows the network to directly learn rotation-aware features from RGB images. As illustrated in Fig. 1, our DVMNet++ requires significantly fewer multiply-accumulate operations (MACs) while achieving smaller angular errors than its hypothesis-based competitors.
@@
We perform comprehensive experiments on the CO3D [23], Objaverse [24], LINEMOD [18], and LINEMOD-O [25] datasets. Our method yields more accurate and robust relative pose estimates for previously unseen objects than the existing competitors. We also conduct ablation studies where the results demonstrate the effectiveness of the key components in our framework. In short, our contributions are threefold:
  • We eliminate the reliance on ground-truth object bounding boxes in relative pose estimation for unseen objects by introducing a new open-set object detector.
  • We tackle the problem of relative object rotation estimation in a hypothesis-free manner by presenting a deep voxel matching network.
  • We present a weighted closest voxel algorithm that robustly computes the relative object rotation from voxel-voxel correspondences in an end-to-end manner.
This paper extends our previous work [26] by highlighting the importance of open-set object detection in the pipeline of relative object pose estimation. We integrate an open-set object detector with our previous DVMNet, achieving generalizable 6D relative object pose estimation without relying on ground-truth object bounding boxes. We also provide a more detailed analysis of our method, showcasing the robustness towards occlusion and the compatibility in the scenario of sparse references.
The remainder of this paper is organized as follows. Section 2 provides an overview of the related work. Section 3 presents the detailed methodology of DVMNet++. Section 4 reports the experimental results across several datasets and includes comprehensive ablation studies. Section 5 summarizes our contributions and outlines directions for future work.
@@ 2 Related Work
Instance-Level Object Pose Estimation. The majority of previous deep learning approaches to object pose estimation [27, 28, 29, 30, 31] tackle the problem at an instance level, assuming that the training and testing data depict the same object instances. Since the appearance of an object instance in different poses typically exhibits limited variations, these methods provide highly accurate object pose estimates. Nevertheless, they struggle to generalize to previously unseen objects during testing without retraining the network, as has been observed in the literature [5, 6, 7]. This limitation constrains their applicability in real-world scenarios that often involve diverse object instances. This problem has been remedied to a degree by category-level object pose estimation methods [32, 33, 34]. In this scenario, the testing images comprise new object instances from specific categories already included in the training data. Although these methods have achieved promising generalization ability within the predefined object categories, they become ineffective when facing objects from entirely new categories.
@@
Generalizable Object Pose Estimation. To tackle the scenario of unseen objects from new categories, there has been growing interest in generalizable object pose estimation. When a textured 3D mesh is available for an unseen object, some approaches [35, 9, 6] suggest generating synthetic images as references by rendering the 3D mesh from various viewpoints. Given a query image that depicts this object, a template matching paradigm is utilized to identify the most similar reference and approximate the object pose in the query as that of the selected reference. Some methods bypass the need for 3D meshes by assuming the availability of multiple real reference images. Object pose estimation is then carried out by employing either a template matching strategy [5] or a 3D object reconstruction technique [7, 10, 8]. Nevertheless, all of these methods rely on having access to dense-view reference images, which limits their applicability in scenarios where only sparse reference views are available.
@@
Relative Object Pose Estimation. In such a context, several studies [12, 13, 14] have highlighted the importance of relative object pose estimation. These methods stand out in generalizable object pose estimation due to their key advantage of requiring only a single reference image. The objective of these methods is to estimate the relative object pose between the input query image and the reference. Since the single-reference assumption tends to result in a large object pose difference between the query and the reference, unseen object pose estimation becomes more challenging. Intuitively, one could establish pixel-pixel correspondences between the two images and compute the relative object pose based on multi-view geometry [15]. However, as reported in the literature [13, 14] and also in our experiments, image-matching techniques [16, 17, 36] have difficulty in delivering accurate pose estimates when confronted with large object pose differences. To address this issue, existing methods [12, 13, 14] suggest approximating the relative object rotation via a discrete set of rotation hypotheses, and learning to maximize the score of the positive hypotheses. Since object rotation lies in a continuous space [37], accurately approximating the rotation necessitates a vast number of rotation hypotheses, which makes such a hypothesis-based approach computationally expensive. Moreover, scoring discrete samples lacks an understanding of the continuous rotation distribution, leading to failure cases with unnaturally high rotation estimation errors. As an alternative, in [38, 39], a diffusion mechanism is employed to regress the pose parameters. The iterative denoising process during inference nonetheless makes this approach time-consuming. By contrast, we present a hypothesis-free technique that is capable of computing the relative object rotation in a single pass via deep voxel matching.
Open-Set Object Detection. The aforementioned relative object pose estimation approaches are designed for object-centric scenarios where the object is positioned at the center of the image and thus the object bounding box is easy to obtain. In this context, the ground-truth object bounding box is assumed to be available. It is employed to predict the relative object translation [12, 13] and crop the object from the query image. However, in some applications, especially in cluttered scenes, effectively detecting a previously unseen object is challenging. In recent years, open-set object detection [19, 20, 21, 22] has received significant attention due to its capacity to identify novel objects from unseen categories. Some pioneering methods have been developed, incorporating open-set object detection into object pose estimation. For instance, CNOS [40] and SAM-6D [41] propose to utilize SAM [42] to generate object proposals. The object mask is selected by matching the proposals with templates based on DINOv2 [43] feature similarities. Gen6D [5] and LocPoseNet [44] predict the bounding box parameters building upon a template matching mechanism. However, a notable limitation of these approaches is their reliance on dense-view reference images, making them inapplicable to our single-reference setting. Therefore, we present a new unseen object detection approach that leverages multi-modal reference information from a single view. As will be witnessed by our experiments, it yields robust detection results for previously unseen objects.
(a) Input
(b) Output
Figure 2: Problem formulation. (a) Input to our method, consisting of a query image and a reference image. (b) Our goal is to identify the corresponding object in the query image and estimate the object translation and rotation based on the reference image. We represent the predicted translation and rotation as a bounding box and green arrows, respectively.
@@ 3 Method
Figure 3: Open-set object detection. We incorporate an open-set object detection module in our relative object pose estimation framework, utilizing multi-modal reference information. Given the reference image, we describe the object appearance using text prompts. An open-vocabulary object detection network takes these prompts and the query image as input, and predicts a set of object proposals. Since the generated proposals may include outliers, we propose identifying the most reliable prediction using an image retrieval technique. We encode the reference image and proposals to feature descriptors by utilizing a pretrained DINOv2 encoder. The final detection result is determined as the proposal with the highest cosine similarity score.
@@ 3.1 Problem Formulation
We tackle the problem of estimating the 6D pose $\mathbf{P}$ for a previously unseen object depicted in an RGB image $\mathbf{I}_{q}$. The object pose consists of a 3D translation $\mathbf{T}$ and a 3D rotation $\mathbf{R}$. In this scenario, the objects present in the testing set $\Omega_{\text{test}}$ differ from those in the training set $\Omega_{\text{train}}$, and the goal is to handle the unseen objects without retraining the network. Furthermore, we assume that one RGB image $\mathbf{I}_{r}$ depicting the object is given as a reference, following the setting in [14, 26]. Notably, both 3D CAD models and dense-view reference images are unavailable in this setting.
Therefore, the goal is to estimate the relative object pose $\Delta{\mathbf{P}}$ between the query image $\mathbf{I}_{q}$ and the reference image $\mathbf{I}_{r}$. The challenges of this problem lie in the ability to generalize to the unseen objects in $\Omega_{\text{test}}$ and in the need for robustness to the large object pose difference between $\mathbf{I}_{q}$ and $\mathbf{I}_{r}$. As illustrated in Fig. 2, unlike previous methods [13, 14, 26], we do not assume a known bounding box to identify the object in the query image. Instead, an open-set object detection approach is required to yield object bounding box parameters $(c_{x},c_{y},w,h)$, where $c_{x}$ and $c_{y}$ denote the center of the bounding box, and $w$ and $y$ indicate the width and height. Furthermore, to estimate the relative pose, we define the reference object coordinate system such that its origin aligns with that of the canonical coordinate system, i.e., setting $\mathbf{T}_{r}={[0,0,0]}^{T}$. The ground-truth relative object translation and rotation are then defined as $\Delta{\mathbf{T}}=\mathbf{T}_{q}$ and $\Delta{\mathbf{R}}=\mathbf{R}_{q}\mathbf{R}_{r}^{T}$, respectively. According to the pinhole camera model, $\mathbf{T}_{q}$ can be computed as
$\mathbf{T}_{q}=d_{q}\mathbf{K}_{q}{[u_{q},v_{q},1]}^{T},$ (1)
where $\mathbf{K}_{q}$ denotes the camera intrinsic, $d_{q}$ represents the depth of the object center, and $(u_{q},v_{q})$ indicates the 2D object center in the query image. In this paper, we assume known camera intrinsics [45] and approximate the 2D object center as the center of the detected bounding box, i.e., $u_{q}=c_{x}$ and $v_{q}=c_{y}$. Since we do not have access to the 3D object model, the translation estimate is inherently ambiguous and can only be determined up to a scale factor. To address the scale ambiguity, we evaluate the translation estimation in terms of the angular error [38]. We will elaborate on this metric in Section 4.
To estimate the relative object rotation, previous hypothesis-based approaches [12, 13, 14] approximate $\Delta{\mathbf{R}}$ by sampling discrete rotation hypotheses and maximizing the score of the positive samples. This can be formulated as
@@
where $\mathcal{R}$ denotes the set of discrete rotation hypotheses. Achieving a decent approximation accuracy requires a large number of hypotheses, e.g., 500,000 in [13]. By contrast, we present a hypothesis-free technique that computes $\Delta{\mathbf{R}}$ in a single pass as $\Delta{\hat{\mathbf{R}}}=g(\mathbf{I}_{q},\mathbf{I}_{r})$.
3.2 Open-Set Object Detection
Recall that existing object detection methods [40, 41, 44] in generalizable object pose estimation rely on dense-view reference images. Consequently, we introduce a new open-set object detection approach that is applicable in our single-reference setting. We propose to facilitate the open-set object detection by leveraging multi-modal reference information, consisting of the RGB image and a natural language description.
Specifically, as shown in Fig. 3, we describe the object in terms of its attributes and category based on the available reference image. The resulting text prompts, along with the query image, are employed as inputs to an open-vocabulary object detector [21]. The output of this detector consists of $M$ object proposals, which lets us formulate the detection process as
$\{p_{1},p_{2},\dots,p_{M}\}=f_{d}(\mathbf{I}_{q},t_{q}|\theta),$ (3)
where $p_{i}$ denotes an object proposal, $f_{d}$ indicates the detection network [21] with pretrained parameters $\theta$, and $t_{q}$ represents the text prompt. Each proposal denotes a region with parameters $(c_{x}^{i},c_{y}^{i},w_{i},h_{i},s_{i})$ in the query image, which is likely to contain the object, with $s_{i}$ a confidence score indicating the reliability of the prediction.
In our initial experiments, we observed the object proposals to be noisy. As illustrated Fig. 3, some detected regions contain the wrong objects. To identify the most reliable result from the candidates, we present an image retrieval strategy utilizing the reference image. We crop the regions from the query image based on the proposal parameters. Each cropped image and the reference image are then encoded into feature descriptors in a high-dimensional space. To this end, we employ a pretrained DINOv2 [43] as the feature encoder and perform spatial average pooling over the output feature map to obtain the descriptor. The cosine similarity between the reference and the proposal is then computed as
$\hat{s}_{i}=\frac{\mathbf{f}_{r}\cdot\mathbf{f}_{i}}{{{||\mathbf{f}_{r}||}_{2}{||\mathbf{f}_{i}||}_{2}}},\ \ \mathbf{f}_{r},\mathbf{f}_{i}\in\mathbb{R}^{C_{d}}$ (4)
where $\mathbf{f}_{r}$ and $\mathbf{f}_{i}$ denote the feature descriptors of the reference and a proposal. We select the proposal with the highest cosine similarity as the final detection result. It is worth noting that one plausible alternative is to utilize the confidence score $s_{i}$ predicted via the open-vocabulary object detector. However, as will be demonstrated in Section 4, such an alternative is less effective than our retrieval-based strategy.
3.3 Hypothesis-Free Relative Rotation Estimation
Given the bounding box predicted by our open-set object detection method, we crop the object from the query image. In this section, we focus on estimating the relative object rotation between the cropped query image and the reference image.
Figure 4: Network architecture of our autoencoder. The encoder takes two RGB images, query and reference, as input and lifts their 2D feature embeddings to 3D voxels by leveraging cross-view 3D information. $\mathbf{O}_{q}$ and $\mathbf{O}_{r}$ represent the learned 3D objectness maps account for robust object rotation estimation. The decoder then reconstructs the masked object images from the voxels, allowing the voxels to encode the object patterns.
3.3.1 Motivation